OpenStudy (anonymous):

)Which of the following expressions can be used to find the area of a square with a side length of fraction 1 over 3 m? fraction 1 over 3 whole squared m2 fraction 1 over 3 whole cubed m2 fraction 1 over 3 multiplied by 2 m2 2 multiplied by fraction 1 over 3 m2

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

um i think its c but im not sure

OpenStudy (anonymous):

area=length*width\[=\frac{ 1 }{ 3 }m ~\times~\frac{ 1 }{ 3 }m=\left( \frac{ 1 }{ 3 } \right)^2~m^2\]

OpenStudy (anonymous):

wait no its A right @surjithayer

OpenStudy (anonymous):

correct